1.Two cars leave at the same time. One car drives North at a rate of 65 m/hr and the other car drives south at a rate of 55 m/hr. In how many hours will the cars be 540 miles apart?

NOTE:( can you please show the work or somehow explain them so i can understand how to do it..thank you)

Found 2 solutions by texttutoring, richwmiller:
Answer by texttutoring(324)

We will make use of the formula D=VT, where D=distance, V=velocity (speed) and T=time
Let Dn,Vn = Distance travelled by the north car, velocity of the north car
Ds, Vs = Distance travelled by the south car, velocity of the south car
The times they travel are the same, because they left at the same time. So tn=ts=t
We know that the two distances, Dn and Ds, add up to 540 miles:
Dn + Ds = 540
Now put in the formulas D=VT, instead of Dn and Ds:
Vn*t + Vs*t = 540
65t + 55t = 540
120t = 540
t = 4.5 hours
The cars will be 540 miles apart in 4.5 hours.

Answer by richwmiller(17219)

r*t=d
speed*time =distance
here we add them both added together to equal 540 miles apart
r*t+r*t=d
65*t+55*t+540
120t=540
t=540/120
t=27/6
t=4 3/6
t=4.5 hours